What is the value of x when 8x-7=5x+2 ?
Answer by nerdybill(7384)

8x-7=5x+2
Add 7 to both sides:
8x-7+7=5x+2+7
8x=5x+9
.
Subtract 5x from both sides:
8x=5x+9
8x-5x=5x+9-5x
3x=9
.
Divide both sides by 3:
(3x)/3=9/3
x=3